# Scalable Learning of Segment-Level Traffic Congestion Functions

### v1

@@ Abstract
We propose and study a data-driven framework for identifying traffic congestion functions (numerical relationships between observations of macroscopic traffic variables) at global scale and segment-level granularity. In contrast to methods that estimate a separate set of parameters for each roadway, ours learns a single black-box function over all roadways in a metropolitan area. First, we pool traffic data from all segments into one dataset, combining static attributes with dynamic time-dependent features. Second, we train a feed-forward neural network on this dataset, which we can then use on any segment in the area. We evaluate how well our framework identifies congestion functions on observed segments and how it generalizes to unobserved segments and predicts segment attributes on a large dataset covering multiple cities worldwide. For identification error on observed segments, our single data-driven congestion function compares favorably to segment-specific model-based functions on highway roads, but has room to improve on arterial roads. For generalization, our approach shows strong performance across cities and road types: both on unobserved segments in the same city and on zero-shot transfer learning between cities. Finally, for predicting segment attributes, we find that our approach can approximate critical densities for individual segments using their static properties.
@@ I Introduction
We consider the problem of identifying segment-level traffic congestion functions, which are numerical relationships between observed values of the macroscopic traffic variables: flow, speed, and density. For any road segment, given potentially partial observations of one variable in a time interval, we want to estimate another (a system identification problem [1]). In particular, our approach learns from data without assumptions on traffic physics and seeks a global solution that can scale to the whole world.
@@ II-A Identifying Fundamental Diagrams
The fundamental diagram (FD) of traffic Haight [6], relates macroscopic traffic variables to each other. A rich history of research exists on proposing and using fundamental diagrams for traffic analysis and management [7].
@@ IV-A Premise and Overview
Empirical flow-speed relationships are noisy and difficult to identify with analytical models [22, 4]. As with other regression tasks, using features (both static segment attributes and time-varying dynamic signals) that plausibly influence the output variable can help improve estimation error. We could have just used features to learn a separate CF for every individual segment. But this approach is also unsatisfying. In principle, if the features influence traffic dynamics, then segments and time-intervals with similar features should have similar flow-to-speed mappings, and we should be able to learn a single CF across segments. Such a CF can both work well for segments with many observations and generalize to roadways with little or no data (which a per-segment CF, whether data-driven or analytical, cannot). In practice, maintaining and updating separate models for every individual roadway in a large urban area could be challenging.
@@ VII Conclusion
Future research could improve several aspects of our framework. More complex segment features (e.g., stop signs or traffic lights) and segment clustering techniques could improve overall accuracy of speed estimates, especially on arterials. The machine learning pipeline could include more inductive biases, such as the graphical structure of the road network, e.g., through graph neural networks, or physics-informed constraints on traffic. We could also solve related but distinct formulations of the overall task, e.g., inferring other traffic properties on the roadways or forecasting congestion in the near future. Finally, evaluating our approach on a large open-source dataset, such as the recently published one from Neun et al. [29] would further prove its value.

### v2

@@ Abstract
We propose and study a data-driven framework for identifying traffic congestion functions (numerical relationships between observations of traffic variables) at global scale and segment-level granularity. In contrast to methods that estimate a separate set of parameters for each roadway, ours learns a single black-box function over all roadways in a metropolitan area. First, we pool traffic data from all segments into one dataset, combining static attributes with dynamic time-dependent features. Second, we train a feed-forward neural network on this dataset, which we can then use on any segment in the area. We evaluate how well our framework identifies congestion functions on observed segments and how it generalizes to unobserved segments and predicts segment attributes on a large dataset covering multiple cities worldwide. For identification error on observed segments, our single data-driven congestion function compares favorably to segment-specific model-based functions on highway roads, but has room to improve on arterial roads. For generalization, our approach shows strong performance across cities and road types: both on unobserved segments in the same city and on zero-shot transfer learning between cities. Finally, for predicting segment attributes, we find that our approach can approximate critical densities for individual segments using their static properties.
@@ I Introduction
We consider the problem of identifying segment-level traffic congestion functions, which are numerical relationships between observed values of the key traffic variables: flow, speed, and density. For any road segment, given potentially partial observations of one variable in a time interval, we want to estimate another (a system identification problem [1]). In particular, our approach learns from data without assumptions on traffic physics and seeks a global solution that can scale to the whole world.
@@ II-A Identifying Fundamental Diagrams
The fundamental diagram (FD) of traffic [6] relates macroscopic traffic variables to each other. A rich history of research exists on proposing and using fundamental diagrams for traffic analysis and management [7].
@@ IV-A Premise and Overview
Empirical flow-speed relationships are noisy and difficult to identify with analytical models [22, 4]. As with other regression tasks, using features (both static segment attributes and time-varying dynamic signals) that plausibly influence the output variable can help improve estimation error. We could have just used features to learn a separate CF for every individual segment, but this approach is also unsatisfying. In principle, if the features influence traffic dynamics, then segments and time-intervals with similar features should have similar flow-to-speed mappings, and we should be able to learn a single CF across segments. Such a CF can both work well for segments with many observations and generalize to roadways with little or no data (which a per-segment CF, whether data-driven or analytical, cannot). In practice, maintaining and updating separate models for every individual roadway in a large urban area could be challenging.
@@ V-C2 Aggregate Error Metrics
A quick comment on the raw loss of predicting inverse speed itself: the mean squared error ranges from $37$ units on highways and $34$ units on arterials.
@@ VII Conclusion
Future research could improve several aspects of our framework. Our segment features and highway/arterial distinctions were simple. More complex features (e.g., stop signs, traffic lights, neighboring segments) and better segment clustering techniques could improve overall estimation accuracy, especially on tricky arterial roads. The machine learning pipeline could include more inductive biases, such as the graphical structure of the road network, e.g., through graph neural networks, or physics-informed constraints.
We could extend the approach to solve related but distinct formulations of the overall task, e.g., inferring other traffic properties on the roadways or forecasting congestion in the near future. Evaluating our approach on a large open-source dataset, such as the one from Neun et al. [29] would further prove its value. Finally, scalable learning-based congestion functions could enable a new approach to citywide mesoscopic simulations that use such models in the loop.
Acknowledgments
The authors would like to thank Pranjal Awasthi, Chun-Ta Lu, and Andrew Tomkins (all from Google Research) for their advice and comments.